OpenStudy (peterg2001):

help wIll medal If n is a term of the sequence 2, 5, 8, 11, ..., which expression would give the value of n? 2n + 1 2n - 1 3n - 1 n + 3 @silverr @mover96 @beastieman21

OpenStudy (peterg2001):

@ichliebedich

OpenStudy (ichliebedich):

(3n-1) does not give the value of n. It gives the value of the nth term of the series. For n=1, the 1st series term is 3*1-1 = 3-1 = 2 For n=2, the 2nd series term is 3*2-1 = 6-1 = 5 For n=3, the 3rd series term is 3*3-1 = 9 -1 = 8 ... (It appears that several of the questions on this quiz are badly worded. This is one of them.) n+3 (answer D